# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@ 3 Riccati-type differentiable coverings determined by ZCRs
Let $\alpha\in\mathfrak{g}\otimes\bar{\Lambda}^{1}\left({\cal E}\right)$ be a ZCR of the form $\alpha=X\,dx+T\,dt$, for a formally integrable equation :=$\mathcal{E}:=\left\{\mathbf{F=0}\right\}$ with two independent variables $\left(x,t\right)$. We assume here that $X$ and $T$ are $\ell\times\ell$ matrices, with $\ell\geq 2$.
@@
This means that, by suitably enlarging the space of dependent variables, $\alpha$ defines a differentiable covering $\tau:\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$, with $\mathcal{Y}:=\left\{\mathbf{F=0},\,\bar{D}_{x}V=XV,\,\bar{D}_{t}V=TV\right\}$. In such an enlarged space, (32) can be also understood as the integrability condition of an equation of the form
@@
Now, it is noteworthy to observe that linear transformations $V\mapsto V^{S}:=SV$, by a nonsingular matrix valued function $S$ on $\mathcal{E}^{(\infty)}$, naturally determine a $\mathcal{C}$-morphism of $\mathcal{Y}$. Under such transformations, (33) transforms as
$V_{x}^{S}=X^{S}V^{S},\qquad V_{t}^{S}=T^{S}V^{S}$
@@
For any fixed $h\in\{1,...,\ell\}$, under the gauge transformation $V\mapsto V^{\prime}=\frac{1}{v^{h}}V$ equation (34) rewrites as
$\left\{\begin{array}[]{l}\bar{d}_{H}\left(\ln v^{h}\right)=\left[\alpha\right]^{h}V^{\prime},\vspace{10pt}\\ \bar{d}_{H}V^{\prime}=\alpha V^{\prime}-\left(\left[\alpha\right]^{h}V^{\prime}\right)V^{\prime}\end{array}\right.$
where $\left[\alpha\right]^{h}$ denotes the $h$-th row of $\alpha$. Moreover, the first order system (33) is equivalent to the following one
$\left\{\begin{array}[]{l}{\displaystyle\frac{v_{x}^{h}}{v^{h}}}=\left[X\right]^{h}V^{\prime},\vspace{10pt}\\ {\displaystyle\frac{v_{t}^{h}}{v^{h}}}=\left[T\right]^{h}V^{\prime},\vspace{10pt}\\ V^{\prime}_{x}=XV^{\prime}-\left(\left[X\right]^{h}V^{\prime}\right)V^{\prime},\\ V^{\prime}_{t}=TV^{\prime}-\left(\left[T\right]^{h}V^{\prime}\right)V^{\prime},\end{array}\right.$
@@
$\rho^{j}:=\left\{\begin{array}[]{ll}v^{j}/v^{h},&j<h,\vspace{10pt}\\ v^{j+1}/v^{h},&j>h.\end{array}\right.$
It follows that the differentiable covering $\tau:\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ is the extension, by means of the potential of a nonlocal conservation law, of a differentiable covering defined by a system of $\ell-1$ Riccati equations.
@@
To any ZCR $\alpha=X\,dx+T\,dt$ of :=$\mathcal{E}:=\left\{\mathbf{F=0}\right\}$, with $\ell\times\ell$ matrices $X$ and $T$, it is associated the $(\ell-1)$-dimensional differentiable covering defined by the system
$\left\{\begin{array}[]{l}\rho_{x}^{j}=\sum_{s=1}^{\ell-1}\left(-X_{\ell s}\rho^{s}\rho^{j}+X_{js}\rho^{s}\right)-X_{\ell\ell}\rho^{j}+X_{j\ell},\qquad j=1,...,\ell-1\vspace{10pt}\\ \rho_{t}^{j}=\sum_{s=1}^{\ell-1}\left(-T_{\ell s}\rho^{s}\rho^{j}+T_{js}\rho^{s}\right)-T_{\ell\ell}\rho^{j}+T_{j\ell},\qquad j=1,...,\ell-1\vspace{10pt}\\ \mathbf{F=0}.\end{array}\right.$
@@
Analogous results hold for any other $h=1,2,...,\ell-1$, with the corresponding nonlocal variables $\bar{\rho}_{1},\bar{\rho}_{2},...,\bar{\rho}_{\ell-1}$ reciprocally related, for any choice of $h$, to the $\rho_{1},\rho_{2},...,\rho_{\ell-1}$ above. However, it is noteworthy to remark that, despite in any case the obtained Riccati systems are equivalent under the transformations $(\rho_{1},\rho_{2},...,\rho_{\ell-1})\mapsto\left(\bar{\rho}_{1},\bar{\rho}_{2},...,\bar{\rho}_{\ell-1}\right)$, the corresponding conservation laws are in general not equivalent.
@@
We consider here KdV equation =$\mathcal{E}=\{z_{xxx}=-z_{t}-6zz_{x}\}$ with the ZCR
$\alpha=\left(\begin{array}[]{cc}0&1\vspace{5pt}\\ -\lambda-z&0\end{array}\right)dx+\left(\begin{array}[]{cc}z_{x}&4\,\lambda-2\,z\vspace{5pt}\\ -4\lambda^{2}-2\,\lambda\,z+2\,z^{2}+z_{xx}&-z_{x}\end{array}\right)dt,\qquad\lambda\in\mathbb{R}.$
In this case, the procedure described above, with $\rho=v^{2}/v^{1}$, provides the $1$-dimensional differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ defined by the system (26) of Example 16, together with the nonlocal conservation law $\mu=\rho dx+\left(z_{x}+\left(4\lambda-2z\right)\rho\right)dt$.
@@
We consider here Tzitzeica equation =$\mathcal{E}=\{z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}}\}$ with the ZCR
@@
$\alpha=\left(\begin{array}[]{ccc}-z_{x}&0&\lambda\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\lambda&z_{x}&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&\lambda&0\end{array}\right)dx+\left(\begin{array}[]{ccc}0&\frac{1}{\lambda}{{\rm e}^{-2\,z}}&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&0&\frac{1}{\lambda}{{\rm e}^{z}}\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\frac{1}{\lambda}{{\rm e}^{z}}&0&0\end{array}\right)dt,\qquad\lambda\in\mathbb{R}.$
@@
In this case, the procedure described above, with $\rho^{1}=v^{1}/v^{3}$ and $\rho^{2}=v^{2}/v^{3}$, provides the $2$-dimensional differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ defined by
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}^{1}=-\lambda\,\rho^{1}\rho^{2}-z_{x}\,\rho^{1}+\lambda,\vspace{5pt}\\ \rho_{t}^{1}=-{{\rm e}^{z}}{\left(\rho^{1}\right)}^{2}+\frac{1}{\lambda}e^{-2z}\rho^{1},\vspace{5pt}\\ \rho_{x}^{2}=-\lambda{\left(\rho^{2}\right)}^{2}+z_{x}{\rho^{2}}+\lambda\rho^{1},\vspace{5pt}\\ \rho_{t}^{2}=-\frac{1}{\lambda}{{\rm e}^{z}}{\rho^{1}}{\rho^{2}}+\frac{1}{\lambda}e^{z},\vspace{5pt}\\ z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}},\end{array}\right.$
together with the nonlocal conservation law
$\mu=\lambda{\small\rho^{2}}\,dx+\frac{1}{\lambda}e^{z}{\small\rho^{1}}\,dt.$
4 Nonlocal pseudo-symmetries and Bäcklund transformations
As discussed in Subsection 2.4, BT are $\mathcal{C}$-morphisms involving differentiable coverings of equations. Our aim here is to illustrate the relationship of BT with nonlocal pseudosymmetries. To this end, let $\mathcal{Y}$ be a system of the form (23) that defines a differentiable covering of $\mathcal{E}$. As discussed in Subsection 2.3.2, assuming that $\mathcal{Y}$ admits a pseudosymmetry $Y$ (or, an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$), for any system of functionally independent invariants $S=\{x^{\prime}_{i}=\xi^{i},\,z^{\prime}{}^{j}=\nu^{j}\}$, satisfying (4), one can consider the corresponding system $\mathcal{Q}$, satisfied by the prolonged invariants $\nu_{\sigma}^{j}$ modulo $\mathcal{E}$ and its consequences. Then, since the form of $\mathcal{Q}$ depends on the choice of the system $S$, one might ask whether it is possible to choose $S$ such that $\mathcal{Q}$ is just a copy of $\mathcal{E}$ (or, another equation $\mathcal{E}^{\prime}$ of interest) or includes it among its differential consequences. With such a choice one would obtain a $\mathcal{C}$-morphism from $\mathcal{Y}$ to $\mathcal{E}$ (or $\mathcal{E}^{\prime}$, respectively), that describes an ABT (or BT, resp.).
@@ 4 Nonlocal pseudo-symmetries and Bäcklund transformations
On the other hand, to calculate pseudosymmetries (or, $r$-pseudosymmetries) of $\mathcal{Y}$, one needs to know its conservation laws $\mu$ (or, horizontal $1$-forms $\gamma$ satisfying the zero-curvature condition) first. Therefore, for the practical application of the procedure described above, it is necessary first to determine not only differential coverings but also some $\mu$ or $\gamma$, should one be interested to $r$-pseudosymmetries. Based on the results of the previous section, it is therefore clear that in the case of an equation with two independent variables ($n=2$) a particularly advantageous situation is that of equations admitting zero-curvature representations: indeed, starting from these representations, the Riccati coverings together with associated nonlocal conservation laws $\mu$ remain automatically determined and one can directly pass to search for corresponding pseudosymmetries. Instead, we do not yet have a result analogous to that of Proposition 23 that can help us construct the horizontal $1$-forms $\gamma$ needed to compute $r$-pseudosymmetries. Hence, calculating the possible $r$-pseudosymmetries admitted by the Riccati coverings, associated with ZCRs, remains rather complicated, as is already the case with the $r$-pseudosymmetries of an equation $\mathcal{E}$, without considering any differentiable covering at all, so much so that for these $r$-pseudosymmetries, it is not easy to find significant examples in the literature. On the other hand, in the context of this work, we think that $r$-pseudosymmetries should play an important role when the Riccati covering has dimension greater than $1$, because in this case the covering equations themselves impose non-trivial compatibility conditions for the invariants of a possible pseudosymmetry, thus indicating the existence in this case of strong restrictions on the possibility of having pseudosymmetries. Therefore, starting from a $\ell\times\ell$ matrix-valued ZCR, with $\ell>2$, looking for $(\ell-1)$-pseudosymmetries is more natural. This is what we did, for instance, in Example 34, where we used a $\mathfrak{sl}_{3}(\mathbb{R})$-valued ZCR of the Tzitzeica equation to show that the ABT found in [9] is also obtainable by following the approach described above, by using a $2$-pseudosymmetry of the corresponding Riccati covering. Thus we believe that the question of the computation of these $r$-pseudosymmetries deserves a further detailed study. The same is true for the more general question on the role that pseudosymmetries and $r$-pseudosymmetries can have in the integrability property of an equation. We also believe that pseudosymmetries can play a fundamental role in studying Bäcklund transformations in higher dimensions. These and other questions will be addressed in future works.

### v5

@@ 3 Riccati-type differentiable coverings determined by ZCRs
In this section, starting from a ZCR of an equation $\mathcal{E}$, we will show how to determine a first-order Riccati-type system defining a differentiable covering together with a corresponding nonlocal conservation law. We will see in Section 4 that this type of covering and the corresponding nonlocal conservation law are fundamental for the computation of pseudosymmetries that allow one to determine Bäcklund transformations.
Let $\alpha\in\mathfrak{g}\otimes\bar{\Lambda}^{1}\left({\cal E}\right)$ be a ZCR of the form $\alpha=X\,dx+T\,dt$, for a formally integrable equation $\mathcal{E}:=\left\{\mathbf{F=0}\right\}$ with two independent variables $\left(x,t\right)$. We assume here that $X$ and $T$ are $\ell\times\ell$ matrices, with $\ell\geq 2$.
@@
This means that, by suitably enlarging the space of dependent variables, $\alpha$ defines a differentiable covering $\mathcal{V}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$, with $\mathcal{V}:=\left\{\mathbf{F=0},\,\bar{D}_{x}V=XV,\,\bar{D}_{t}V=TV\right\}$. In such an enlarged space, (36) can be also understood as the integrability condition of an equation of the form
@@
Now, it is noteworthy to observe that linear transformations $V\mapsto V^{S}:=SV$, by a nonsingular matrix valued function $S$ on $\mathcal{E}^{(\infty)}$, naturally determine a $\mathcal{C}$-morphism of $\mathcal{V}$. Under such transformations, (37) transforms as
$V_{x}^{S}=X^{S}V^{S},\qquad V_{t}^{S}=T^{S}V^{S},$
@@
For any fixed $h\in\{1,$$\ell\}$, under the gauge transformation $V\mapsto V^{\prime}=\frac{1}{v^{h}}V$ equation (38) rewrites as
$\left\{\begin{array}[]{l}\bar{d}_{H}\left(\ln v^{h}\right)=\left[\alpha\right]^{h}V^{\prime},\vskip 10.0pt\\ \bar{d}_{H}V^{\prime}=\alpha V^{\prime}-\left(\left[\alpha\right]^{h}V^{\prime}\right)V^{\prime},\end{array}\right$. (39)
where $\left[\alpha\right]^{h}$ denotes the $h$-th row of $\alpha$. Thus, the first order system (37) is equivalent to the following one
$\left\{\begin{array}[]{l}{\displaystyle\frac{v_{x}^{h}}{v^{h}}}=\left[X\right]^{h}V^{\prime},\vskip 10.0pt\\ {\displaystyle\frac{v_{t}^{h}}{v^{h}}}=\left[T\right]^{h}V^{\prime},\vskip 10.0pt\\ V^{\prime}_{x}=XV^{\prime}-\left(\left[X\right]^{h}V^{\prime}\right)V^{\prime},\\ V^{\prime}_{t}=TV^{\prime}-\left(\left[T\right]^{h}V^{\prime}\right)V^{\prime},\end{array}\right$. (40)
@@
$\rho^{j}:=\left\{\begin{array}[]{ll}v^{j}/v^{h},&j<h,\vskip 10.0pt\\ v^{j+1}/v^{h},&j>h.\end{array}\right$.
It follows that the differentiable covering $\mathcal{V}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ is the extension, by means of the potential of a nonlocal conservation law, of a differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$defined by a system of $\ell-1$ Riccati-type equations. This generalises a fact already discovered by Chern and Tenenblat in [15], in the case of $\mathfrak{sl}_{2}(\mathbb{R})$-valued ZCRs for equations describing pseudospherical surfaces.
@@
To any zero-curvature representation $\alpha=X\,dx+T\,dt$ of $\mathcal{E}:=\left\{\mathbf{F=0}\right\}$, with $\ell\times\ell$ matrices $X$ and $T$, it is associated the $(\ell-1)$-dimensional differentiable covering defined by the system
$\left\{\begin{array}[]{l}\rho_{x}^{j}={\displaystyle\sum_{s=1}^{\ell-1}\left(-X_{\ell s}\rho^{s}\rho^{j}+X_{js}\rho^{s}\right)-X_{\ell\ell}\rho^{j}+X_{j\ell},\qquad j=1,...,\ell-1}\vskip 10.0pt\\ \rho_{t}^{j}={\displaystyle\sum_{s=1}^{\ell-1}\left(-T_{\ell s}\rho^{s}\rho^{j}+T_{js}\rho^{s}\right)-T_{\ell\ell}\rho^{j}+T_{j\ell},\qquad j=1,...,\ell-1}\vskip 10.0pt\\ \mathbf{F=0}.\end{array}\right$.
@@
Analogous results hold for any other $h=1,2,...,\ell-1$, with the corresponding nonlocal variables $\bar{\rho}_{1},\bar{\rho}_{2},...,\bar{\rho}_{\ell-1}$ reciprocally related, for any choice of $h$, to the $\rho_{1},\rho_{2},...,\rho_{\ell-1}$ above.
@@
We consider here KdV equation $\mathcal{E}=\{z_{xxx}=-z_{t}-6zz_{x}\}$ with the ZCR
$\alpha=\left(\begin{array}[]{cc}0&1\vskip 5.0pt\\ -\lambda-z&0\end{array}\right)dx+\left(\begin{array}[]{cc}z_{x}&4\,\lambda-2\,z\vskip 5.0pt\\ -4\lambda^{2}-2\,\lambda\,z+2\,z^{2}+z_{xx}&-z_{x}\end{array}\right)dt,\qquad\lambda\in\mathbb{R}.$
In this case, the procedure described above, with $\rho=v^{2}/v^{1}$, provides the $1$-dimensional differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ defined by the system (27) of Example 16, together with the nonlocal conservation law $\mu=\rho dx+\left(z_{x}+\left(4\lambda-2z\right)\rho\right)dt$. According to first equation of (39) $\mu=\bar{d}_{H}(\ln v^{1})$, thus $\rho=v_{x}^{1}/v^{1}$.
@@
We consider here Tzitzeica equation $\mathcal{E}=\{z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}}\}$ with the ZCR
@@
$\alpha=\left(\begin{array}[]{ccc}-z_{x}&0&\lambda\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\lambda&z_{x}&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&\lambda&0\end{array}\right)dx+\left(\begin{array}[]{ccc}0&{\displaystyle\frac{1}{\lambda}{{\rm e}^{-2\,z}}}&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&0&{\displaystyle\frac{1}{\lambda}{{\rm e}^{z}}}\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr{\displaystyle\frac{1}{\lambda}{{\rm e}^{z}}}&0&0\end{array}\right)dt,\qquad\lambda\in\mathbb{R}.$
@@
In this case, the procedure described above, with $\rho^{1}=v^{1}/v^{3}$ and $\rho^{2}=v^{2}/v^{3}$, provides the $2$-dimensional differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ defined by the system (30) of Example 17, together with the nonlocal conservation law $\mu=\lambda{\small\rho^{2}}\,dx+e^{z}{\small\rho^{1}/\lambda}\,dt$. In this case, according to first equation of (39) $\mu=\bar{d}_{H}(\ln v^{3})$, thus $\rho^{1}=\lambda e^{-z}v_{t}^{3}/v^{3}$ and $\rho^{2}=\lambda^{-1}v_{x}^{3}/v^{3}$.
4 Nonlocal pseudosymmetries and Bäcklund transformations
As discussed in Subsection 2.4, a Bäcklund transformation from $\mathcal{E}$ to $\mathcal{E}^{\prime}$ can be seen as a nonlocal $\mathcal{C}$-morphism, from a differentiable extension $\mathcal{V}$ of the equation $\mathcal{E}$ to the equation $\mathcal{E}^{\prime}$. Our aim here is to show how such transformations can result from factorisations of differentiable extensions $\mathcal{V}$ with respect to nonlocal pseudosymmetries. In these cases one specialises to $\mathcal{V}$ the factorisation procedure described in the end of Subsection 2.3.2.
The basic assumption is that $\mathcal{E}$ admits a differentiable extension $\mathcal{V}$ which has a pseudosymmetry $\bar{Y}$ (or an $r$-pseudosymmetry $\{\bar{Y}_{1},...,\bar{Y}_{r}\}$) which leads to a quotient system $\mathcal{Q}$ that, among its equations, has a copy of equation $\mathcal{E}^{\prime}$. In such a case, the Bäcklund transformation is determined by a suitable choice of some basic system of functionally independent invariants $S=\{x^{\prime}_{i}=\xi^{i},\,z^{\prime}{}^{j}=\nu^{j}\}$ of the exploited pseudosymmetry (resp., $r$-pseudosymmetry). Indeed, since the concrete form of $\mathcal{Q}$ depends on the particular choice of $S$, if one chooses another basic system of functionally independent invariants $\hat{S}=\{\hat{x}^{\prime}_{i}=\hat{\xi}^{i},\,\hat{z}^{\prime}{}^{j}=\hat{\nu}^{j}\}$ equivalent to $\{\xi^{i},\,\nu^{j}\}$, i.e., such that $\{\hat{\xi}^{i}=A^{i}(\xi^{i},\nu^{j}),\,\hat{\nu}^{j}=B^{j}(\xi^{i},\nu^{j})\}$ and $\det\left(\tilde{D}_{s}\hat{\xi}^{i}\right)\neq 0$ (where $\tilde{D}_{s}$ are the total derivatives on $\mathcal{V}^{(\infty)}$), one would obtain a system $\mathcal{\hat{Q}}$ which is point equivalent to $\mathcal{Q}$, through the point transformation $\{\hat{x}^{\prime}_{i}=A^{i}(x^{\prime},z^{\prime}),\,\hat{z}^{\prime}{}^{j}=B^{j}(x^{\prime},z^{\prime})\}$.
On the other hand, in view of Proposition 11 and Proposition 12, the infinitesimal conditions (18) and (23) for pseudosymmetries and $r$-pseudosymmetries are linear only with respect to $\varphi$ and $\Phi$. Thus, in general the determination of pseudosymmetries and $r$-pseudosymmetries of $\mathcal{V}$ is feasible if one already knows some $\mu$ and $\gamma$. It follows that, for the practical application of the procedure described above, it may be necessary to already have available the possible differential coverings $\mathcal{V}$ together with some corresponding $1$-forms $\mu$ and $\gamma$, defined in $\mathcal{V}$.
@@ 4 Nonlocal pseudosymmetries and Bäcklund transformations
Based on the results of the previous section, it is therefore clear that in the case of an equation with two independent variables ($n=2$) a particularly advantageous situation is that of equations admitting ZCRs: indeed, starting from these representations, the Riccati-type differentiable coverings together with associated horizontally closed $1$-forms $\mu$ (that are conservation laws when $n=2$) remain automatically determined and one can directly pass to search for corresponding pseudosymmetries. Unfortunately, one does not have yet an analogous result for $r$-pseudosymmetries. Hence, calculating the possible $r$-pseudosymmetries admitted by the Riccati-type differentiable coverings, associated with ZCRs, remains rather complicated; as is already the case without considering any differentiable covering so much so that in the literature it is not easy to find significant examples of $r$-pseudosymmetries of differential equations. Yet we believe that r-pseudosymmetries should play an important role when the Riccati-type differentiable covering has dimension greater than $1$, because the larger number of covering equations in this case may lead to greater restrictions on the existence of pseudosymmetries; therefore, in the case of $\ell\times\ell$ matrix-valued ZCR, with $\ell>2$, the more general $r$-pseudosymmetries should play a key role. For instance, in Example 35, we show that the auto-Bäcklund transformation of the Tzitzeica equation found in [9] can be obtained by using a $2$-pseudosymmetry of the covering obtained from an $\mathfrak{sl}_{3}(\mathbb{R})$-valued ZCR. Thus, the computation of $r$-pseudosymmetries deserves further detailed study. The same is true for the more general question on the role that pseudosymmetries and $r$-pseudosymmetries can have in the integrability property of an equation; in particular, we believe that they can play a fundamental role in studying Bäcklund transformations in higher dimensions. These and other questions will be addressed in future works.